# Test-time Adaptation for Regression by Subspace Alignment

### v1

@@ 4 Experiment
This section provides empirical analysis of feature subspaces and experimental evaluations of SSA on various regression tasks. First, we checked whether the learned features are distributed in a small subspace (Sec. 4.3.1) and then evaluated the regression performance (Secs. 4.3.2 and 4.3.3). We also analyzed the effect of the TTA methods from the perspective of the feature subspace (Sec. 4.3.4).
@@ 4.1 Dataset
SVHN-MNIST. SVHN [44] and MNIST [36] are famous digit-recognition datasets. Although they are mainly used for classification, we used them for regression by training models to directly output a scalar value of the label. We used SVHN and MNIST as the source and target domains, respectively.
UTKFace [59]. UTKFace is a dataset consisting of face images. The task is to predict the age of the person in an input image. For the source model, we trained models on the original UTKFace images. For the target domain, we added corruptions such as noise or blur to the images. The types of corruption were the same as those of ImageNet-C [25]. We applied 13 types of corruption at the highest severity level of the five levels.
@@
California Housing [47]. California Housing is a tabular dataset aiming at predicting housing prices from the information of areas. We split the dataset into non-coastal and coastal areas for the source and target domains in accordance with He et al. [24].
More details of the datasets are provided in Sec. C.1.
@@ 4.2 Setting
Source model. We used ResNet-26 [23] for SVHN, ResNet-50 for UTKFace and Biwi Kinect, and an MLP for California Housing. We modified the last fully-connected layer to output single scalar values and trained the models with the standard MSE loss on each dataset and task. The details of the training are provided in Appendix C.
Test-time adaptation with SSA (ours). We minimized $\mathcal{L}_{\text{TTA}}$ on the target datasets. We used the outputs of the penultimate layer of the model as features, which had 2048 dimensions. We set the number of dimensions of the feature subspace to $K=100$ as the default throughout the experiments. More detailed settings are provided in Sec. C.3.
Baseline. Since there are no TTA baselines designed for regression, we compared SSA with TTA methods designed for classification but can be naively modified to regression: Source (no adaptation), BN-adapt [4], Feature restoration (FR) [16], Prototype, Variance minimization (VM), and RSD [6]. In addition, we used the following methods other than TTA as baselines: test-time training (TTT) [53], DANN [21], and Oracle (fine-tuning using labels; performance upper bound). The details of the baseline methods are described in Sec. C.3.
4.3 Results
@@ 4.3.1 Number of Dimensions of the Feature Subspace
After the pre-training on the source dataset, we counted the numbers of valid feature dimensions (i.e., having non-zero variances) and dimensions of the feature subspace in which the source features are distributed. The latter value corresponds to the rank of the covariance matrix of the source features in Eq. 5. Tab. 1 shows the result of each source dataset and regression test $R^{2}$ scores. Although the number of feature dimensions is 2048 in ResNet, many feature dimensions of the regression models have zero variance because of ReLU activation. This is the cause of the failure of the naive feature alignment, as described in Sec. 3.1. Moreover, the source features are distributed in only a small subspace with fewer than a hundred dimensions. In the California Housing dataset, we can also see the same tendency that the number of the subspace dimensions is only 40 whereas the number of the entire feature dimensions is 100. This property limits the performance of the naive feature alignment in regression since aligning the entire feature space is ineffective to the subspace in which the features are actually distributed.
@@ 4.3.2 Regression Performance
Table 2: Test $R^{2}$ score (higher is better) and RMSE (lower is better) on SVHN-MNIST. The best scores are bolded.
Method $R^{2}~{}(\uparrow)$ RMSE ($\downarrow$)
Source 0.4060.406{0.406}0.406 2.2322.232{2.232}2.232
DANN ${0.307_{\pm 0.09}}$ ${2.406_{\pm 0.16}}$
TTT ${0.288_{\pm 0.02}}$ ${2.443_{\pm 0.03}}$
BN-adapt ${0.396_{\pm 0.00}}$ ${2.251_{\pm 0.01}}$
Prototype ${0.491_{\pm 0.00}}$ ${2.065_{\pm 0.01}}$
FR ${0.369_{\pm 0.01}}$ ${2.300_{\pm 0.02}}$
VM ${-685_{\pm 27.6}}$ ${75.83_{\pm 1.52}}$
RSD ${0.312_{\pm 0.08}}$ ${2.397_{\pm 0.15}}$
SSA (ours) $\mathbf{0.511_{\pm 0.03}}$ $\mathbf{2.024_{\pm 0.06}}$
Oracle ${0.874_{\pm 0.00}}$ ${1.028_{\pm 0.00}}$
@@
Method $R^{2}~{}(\uparrow)$ RMSE ($\downarrow$)
Source 0.6050.605{0.605}0.605 0.6840.684{0.684}0.684
BN-adapt ${0.318_{\pm 0.00}}$ ${0.899_{\pm 0.00}}$
Prototype ${-0.726_{\pm 0.01}}$ ${1.431_{\pm 0.00}}$
FR ${0.510_{\pm 0.01}}$ ${0.762_{\pm 0.01}}$
RSD - -
SSA (ours) $\mathbf{0.639_{\pm 0.00}}$ $\mathbf{0.655_{\pm 0.00}}$
Oracle ${0.729_{\pm 0.00}}$ ${0.567_{\pm 0.00}}$
We evaluated the regression performance in terms of the $R^{2}$ score (coefficient of determination), which is widely used in regression tasks (see Appendix B). Tabs. 2 and 3 show the scores for the SVHN-MNIST and California Housing. In the both cases, SSA outperformed the baselines; some of them even underperformed the Source. This is because the baselines were designed for classification tasks and they broke the feature subspaces learned by the source model (see Sec. 4.3.4). On the other hand, RSD [6] is originally designed for regression in UDA but did not work in the California Housing dataset because of numerical instability of SVD performed on every target feature batch. In contrast, our method is stable since it avoids degenerated dimensions during TTA. Tab. 4 shows the $R^{2}$ scores on the UTKFace with image corruption. We can see that SSA had the highest $R^{2}$ scores for most of the corruption types. In particular, SSA outperformed the baselines by a large margin on noise-type corruption which significantly degraded the performance of Source. Tab. 5 shows the $R^{2}$ scores on Biwi Kinect with genders different from the source domains. SSA constantly had higher $R^{2}$ scores than the baselines; the baselines’ scores sometimes significantly dropped or even diverged (Prototype). In summary, SSA consistently improved the scores whereas the baselines sometimes even underperformed Source. Moreover, SSA worked well not only on image data but also tabular data.
Table 4: Test $R^{2}$ scores on UTKFace with image corruption. The best scores are bolded.
@@
Table 5: Test $R^{2}$ scores on Biwi Kinect. The best scores are bolded.
@@ 4.3.3 Ablation Study
We investigated the effect of subspace detection and dimension weighting by running SSA with and without them. For the SSA variant without subspace detection (i.e., naively aligning the entire feature space), we simply selected the top-$K$ feature dimensions that had the largest variances. In this case, we directly used the weight of the linear regressor $h_{\psi}$ to compute the dimension weight $\alpha_{d}$ as $\alpha_{d}=1+|w_{d}|$ instead of Eq. 6. Tab. 6 shows the test $R^{2}$ scores with and without subspace detection and dimension weighting on each dataset. Without subspace detection, the scores were worse than Source on MNIST and Biwi Kinect, and of the same level as simple baselines like BN-adapt [4] on UTKFace (Tab. 4). In contrast, subspace detection significantly improved the scores on all three datasets. Dimension weighting also improved the scores, although the gain was smaller than in the case of subspace detection. This is because the variance of the feature subspace dimension correlates with the weight; i.e., the top-$K$ selected dimensions with respect to variance tended to have high importance to the output. Tab. 7 lists the correlation coefficients between the top $K=100$ variances of the source features along the source bases $\lambda_{d}^{\text{s}}$ and the corresponding dimension weight $\alpha_{d}$. We can see that there are strong correlations in the three datasets we used. But we can expect that dimension weighting can raise the importance of the feature dimensions that have low variance but affect the output, which further improves regression performance.
Table 7: Correlation coefficients between the top $K=100$ variances of source features along the source bases $\lambda^{\text{s}}_{d}$ and weight $\alpha_{d}$ in Eq. 6.
@@ 4.3.4 Feature Subspace Analysis
Feature reconstruction. To verify that the reason why the baseline methods degrade the regression performance is that they affect the feature subspace learned by the source model as mentioned in Sec. 4.3.2, we examined the reproducibility of the target features with the source bases $\mathbf{V}^{\text{s}}$ after TTA. That is, the target features can be represented by a linear combination of the source bases if the model retains the source subspace throughout TTA and the target features fit within the subspace. To measure this quantitatively, we computed the reconstruction error $L$ as the Euclidean distance between a target feature vector $\mathbf{z}^{\text{t}}$ and $\mathbf{z}_{\text{r}}^{\text{t}}$, the one reconstructed with $n$ source bases:
$L$ $=\|\mathbf{z}_{\text{r}}^{\text{t}}-\mathbf{z}^{\text{t}}\|_{2},$ (9)
$\mathbf{z}_{\text{r}}^{\text{t}}$ $=\bm{\mu}^{\text{s}}+\sum_{d=1}^{n}((\mathbf{z}^{\text{t}}-\bm{\mu}^{\text{s}})^{\top}\mathbf{v}^{\text{s}}_{d})\mathbf{v}^{\text{s}}_{d},$ (10)
where $\mathbf{z}^{\text{t}}$ is a target feature vector extracted with the model after TTA, and $n$ is the number of dimensions of the source subspace listed in Tab. 1.
@@
Figure 2: Reconstruction error of features reconstructed with the source bases relative to the original target features. Note that Source and Prototype are the same, since Prototype does not update the feature extractor of the model. FR [16] and VM are not plotted in (a) because they had huge errors.
Fig. 2 plots the reconstruction error $L$ versus $n$ on the three datasets. The error decreased as $n$ increased for all methods, but SSA reduced the error with a smaller $n$ than in those of the baselines, indicating that it could make the target features fit within the source subspace. Especially in the case of Biwi Kinect (c), the baseline methods produced larger errors than Source; i.e., they broke the learned subspace.
In terms of feature alignment, Fig. 4 in Sec. D.2 visualizes and evaluates the feature gap between the domains in the subspace.
Another effect of subspace detection. Subspace detection has another effect on feature alignment. We used the KL divergence between two diagonal Gaussian distributions in Eq. 4 to measure the distribution gap between the source and target features, under the assumption that the features follow a Gaussian distribution. Here, while it is not clear that the assumption actually holds especially when features are output through activation functions like ReLU as in ResNet, we argue that the subspace detection of SSA has, in addition to an effective feature alignment, the effect of making such features follow a distribution close to a Gaussian.
@@
Figure 3: Histograms of three randomly selected target feature dimension. Left: Original features. Right: Projected features.
We visualized the histograms of the target features $\mathbf{z}_{i}^{\text{t}}$ extracted with the source model and ones projected to the source subspace with Eq. 7. Fig. 3 show the histograms of three randomly selected dimensions of the original and projected features of the three datasets. In the left column of each figure, the histograms of the original features concentrate on zero because of the ReLU activation and do not follow a Gaussian distribution, which makes the KL divergence computation with Eq. 4 inaccurate. On the other hand, the histograms of the projected features in the right columns are close to Gaussians. Thus, subspace detection makes it easier to align the features with the Gaussian KL divergence.
The reason why the projected features follow a Gaussian distribution can be interpreted as follows. From Eq. 7, the $k$-th element of a projected feature vector $\tilde{z}_{i,k}^{\text{t}}$ is
@@ 5.1 Unsupervised Domain Adaptation
Unsupervised domain adaptation (UDA) has been actively studied as a way to transfer knowledge in the source domain to the target domain [9]. Theoretically, it is known that the upper bound of the error on the target domain includes a distribution gap term between the source and target domains [3, 21, 45]. For regression, Cortes and Mohri [8] theoretically explored regression UDA. RSD [6] and DARE-GRAM [43] take into account that the feature scale matters in regression and explicitly align the feature scale during the feature alignment. However, UDA requires the source and target datasets to be accessed simultaneously during training, which can be restrictive when datasets cannot be accessed due to privacy or security concerns, or storage limitations.
More recently, source-free domain adaptation (SFDA), which does not access the source dataset during adaptation, has been studied. The SFDA setting is similar to TTA in that SFDA adapts models with only unlabeled target data. However, SFDA requires to store the whole target dataset and access the dataset for multiple epochs to train additional models [37, 56, 7, 51, 24] or perform clustering [39]. On the other hand, TTA does not train additional models nor access the target dataset for multiple epochs, which enables instant adaptation with low computational resource and storage.

### v2

@@ 4 Experiment
This section provides empirical analysis of feature subspaces and evaluations of SSA on various regression tasks. First, we checked whether the learned features are distributed in a small subspace (Section 4.3.1) and then evaluated the regression performance (Sections 4.3.2 and 4.3.3). We also analyzed the effect of the TTA methods from the perspective of the feature subspace (Section 4.3.4).
@@ 4.1 Dataset
SVHN-MNIST. SVHN (Netzer et al., 2011) and MNIST (LeCun et al., 1998b) are famous digit-recognition datasets. Although they are mainly used for classification, we used them for regression by training models to directly output a scalar value of the label. We used SVHN and MNIST as the source and target domains, respectively.
UTKFace (Zhang et al., 2017). UTKFace is a dataset consisting of face images. The task is to predict the age of the person in an input image. For the source model, we trained models on the original UTKFace images. For the target domain, we added corruptions such as noise or blur to the images. The types of corruption were the same as those of ImageNet-C (Hendrycks & Dietterich, 2019). We applied 13 types of corruption at the highest severity level of the five levels.
@@
California Housing (Nugent, 2017). California Housing is a tabular dataset aiming at predicting housing prices from the information of areas. We split the dataset into non-coastal and coastal areas for the source and target domains in accordance with He et al. (2024).
More details of the datasets are provided in Section C.1.
@@ 4.2 Setting
Source model. We used ResNet-26 (He et al., 2016) for SVHN, ResNet-50 for UTKFace and Biwi Kinect, and an MLP for California Housing. We modified the last fully-connected layer to output single scalar values and trained the models with the standard MSE loss on each dataset and task. The details of the training are provided in Appendix C.
Test-time adaptation with SSA (ours). We minimized $\mathcal{L}_{\text{TTA}}$ on the target datasets. We used the outputs of the penultimate layer of the model as features, which had 2048 dimensions. We set the number of dimensions of the feature subspace to $K=100$ as the default throughout the experiments. More detailed settings are provided in Section C.3.
Baseline. Since there are no TTA baselines designed for regression, we compared SSA with TTA methods designed for classification but can be naively modified to regression: Source (no adaptation), BN-adapt (Benz et al., 2021), Feature restoration (FR) (Eastwood et al., 2022), Prototype, Variance minimization (VM), and RSD (Chen et al., 2021). In addition, we used the following methods other than TTA as baselines: test-time training (TTT) (Sun et al., 2020), DANN (Ganin et al., 2016), and Oracle (fine-tuning using labels; performance upper bound). The details of the baseline methods are described in Section C.3.
4.3 Result
@@ 4.3.1 Number of Dimensions of the Feature Subspace
After the pre-training on the source dataset, we counted the numbers of valid feature dimensions (i.e., having non-zero variances) and dimensions of the feature subspace in which the source features are distributed. The latter value corresponds to the rank of the covariance matrix of the source features in Equation 4. Table 1 shows the result of each source dataset and regression test $R^{2}$ scores. Although the number of feature dimensions is 2048 in ResNet, many feature dimensions of the regression models have zero variance because of ReLU activation. This is the cause of the failure of the naive feature alignment, as described in Section 3.1. Moreover, the source features are distributed in only a small subspace with fewer than a hundred dimensions. In the California Housing dataset, we can also see the same tendency that the number of the subspace dimensions is only 40 whereas the number of the entire feature dimensions is 100. This property limits the performance of the naive feature alignment in regression since aligning the entire feature space is ineffective to the subspace in which the features are actually distributed.
In the MLP used for the California Housing dataset, the subspace dimensions compared to the entire feature space is higher than the ResNets used for the other datasets. One explanation for this difference is model capacity and task complexity. MLP’s capacity is low relative to the task (California Housing)’s complexity. On the other hand, the ResNet-26, which has high capacity, resulted in lower subspace dimensions on SVHN, which has low complexity.
@@ 4.3.2 Regression Performance
Table 2: Test scores on SVHN-MNIST. The best scores are bolded.
Method $R^{2}(\uparrow)$ RMSE ($\downarrow$) MAE ($\downarrow$)
Source 0.4060.406{0.406}0.406 2.2322.232{2.232}2.232 1.6081.608{1.608}1.608
DANN ${0.307_{\pm 0.09}}$ ${2.406_{\pm 0.16}}$ ${1.489_{\pm 0.09}}$
TTT ${0.288_{\pm 0.02}}$ ${2.443_{\pm 0.03}}$ ${1.597_{\pm 0.03}}$
BN-adapt ${0.396_{\pm 0.00}}$ ${2.251_{\pm 0.01}}$ ${1.458_{\pm 0.00}}$
Prototype ${0.491_{\pm 0.00}}$ ${2.065_{\pm 0.01}}$ ${1.479_{\pm 0.01}}$
FR ${0.369_{\pm 0.01}}$ ${2.300_{\pm 0.02}}$ ${1.631_{\pm 0.02}}$
VM ${-685.1_{\pm 27.63}}$ ${75.83_{\pm 1.52}}$ ${75.78_{\pm 1.52}}$
RSD ${0.252_{\pm 0.12}}$ ${2.497_{\pm 0.20}}$ ${1.703_{\pm 0.20}}$
SSA (ours) $\mathbf{0.511_{\pm 0.03}}$ $\mathbf{2.024_{\pm 0.06}}$ $\mathbf{1.209_{\pm 0.04}}$
Oracle ${0.874_{\pm 0.00}}$ ${1.028_{\pm 0.00}}$ ${0.575_{\pm 0.00}}$
@@
Method $R^{2}(\uparrow)$ RMSE ($\downarrow$) MAE ($\downarrow$)
Source 0.6050.605{0.605}0.605 0.6840.684{0.684}0.684 0.5160.516{0.516}0.516
BN-adapt ${0.318_{\pm 0.00}}$ ${0.899_{\pm 0.00}}$ ${0.699_{\pm 0.00}}$
Prototype ${-0.726_{\pm 0.01}}$ ${1.431_{\pm 0.00}}$ ${1.196_{\pm 0.00}}$
FR ${0.510_{\pm 0.01}}$ ${0.762_{\pm 0.01}}$ ${0.534_{\pm 0.01}}$
RSD - - -
SSA (ours) $\mathbf{0.639_{\pm 0.00}}$ $\mathbf{0.655_{\pm 0.00}}$ $\mathbf{0.469_{\pm 0.00}}$
Oracle ${0.729_{\pm 0.00}}$ ${0.567_{\pm 0.00}}$ ${0.404_{\pm 0.00}}$
We evaluated the regression performance in terms of the $R^{2}$ score (coefficient of determination), which is widely used in regression tasks (see Appendix B). Tables 3 and 3 show the scores for the SVHN-MNIST and California Housing. In the both cases, SSA outperformed the baselines; some of them even underperformed the Source. This is because the baselines were designed for classification tasks and they broke the feature subspaces learned by the source model (see Section 4.3.4). On the other hand, RSD (Chen et al., 2021) is originally designed for regression in UDA but did not work in the California Housing dataset because of numerical instability of SVD performed on every target feature batch. In contrast, our method is stable since it avoids degenerated dimensions during TTA. Table 4 shows the $R^{2}$ scores on the UTKFace with image corruption. We can see that SSA had the highest $R^{2}$ scores for most of the corruption types. In particular, SSA outperformed the baselines by a large margin on noise-type corruption which significantly degraded the performance of Source. Table 5 shows the $R^{2}$ scores on Biwi Kinect with genders different from the source domains. SSA constantly had higher $R^{2}$ scores than the baselines; the baselines’ scores sometimes significantly dropped or even diverged (Prototype). In summary, SSA consistently improved the scores whereas the baselines sometimes even underperformed Source. Moreover, SSA worked well not only on image data but also tabular data.
Table 4: Test $R^{2}$ scores on UTKFace with image corruption (higher is better). The best scores are bolded.
@@
Table 5: Test $R^{2}$ scores on Biwi Kinect (higher is better). The best scores are bolded.
@@ 4.3.3 Ablation Study
We performed an ablation study on the subspace detection and dimension weighting. For the SSA variant without subspace detection (i.e., naively aligning the entire feature space), we simply selected the top-$K$ feature dimensions that had the largest variances. In this case, we directly used the weight of the linear regressor $h_{\psi}$ to compute the dimension weight $\alpha_{d}$ as $\alpha_{d}=1+|w_{d}|$ instead of Equation 5. Table 6 shows the test $R^{2}$ scores with and without subspace detection and dimension weighting on each dataset. Without subspace detection, the scores were worse than Source on MNIST and Biwi Kinect, and of the same level as simple baselines like BN-adapt (Benz et al., 2021) on UTKFace (Table 4). In contrast, subspace detection significantly improved the scores on all three datasets. Dimension weighting also improved the scores, although the gain was smaller than in the case of subspace detection. This is because the variance of the feature subspace dimension correlates with the weight; i.e., the top-$K$ selected dimensions with respect to variance tended to have high importance to the output. Table 7 lists the correlation coefficients between the top $K=100$ variances of the source features along the source bases $\lambda_{d}^{\text{s}}$ and the corresponding dimension weight $\alpha_{d}$. We can see that there are strong correlations in the three datasets we used. But we can expect that dimension weighting can raise the importance of the feature dimensions that have low variance but affect the output, which further improves regression performance.
Table 7: Correlation coefficients between the top $K=100$ variances of source features along the source bases $\lambda^{\text{s}}_{d}$ and weight $\alpha_{d}$ in Equation 5.
@@ 4.3.4 Feature Subspace Analysis
Feature reconstruction. To verify that the reason why the baseline methods degrade the regression performance is that they affect the feature subspace learned by the source model as mentioned in Section 4.3.2, we examined the reproducibility of the target features with the source bases $\mathbf{V}^{\text{s}}$ after TTA. That is, the target features can be represented by a linear combination of the source bases if the model retains the source subspace throughout TTA and the target features fit within the subspace. To measure this quantitatively, we computed the reconstruction error $L$ as the Euclidean distance between a target feature vector $\mathbf{z}^{\text{t}}$ and $\mathbf{z}_{\text{r}}^{\text{t}}$, the one reconstructed with $n$ source bases:
$L=\|\mathbf{z}_{\text{r}}^{\text{t}}-\mathbf{z}^{\text{t}}\|_{2},\quad\mathbf{z}_{\text{r}}^{\text{t}}=\bm{\mu}^{\text{s}}+\sum_{d=1}^{n}((\mathbf{z}^{\text{t}}-\bm{\mu}^{\text{s}})^{\top}\mathbf{v}^{\text{s}}_{d})\mathbf{v}^{\text{s}}_{d},$ (8)
where $\mathbf{z}^{\text{t}}$ is a target feature vector extracted with the model after TTA, and $n$ is the number of dimensions of the source subspace listed in Table 1.
@@
Figure 2: Reconstruction error of features reconstructed with the source bases relative to the original target features. Note that Source and Prototype are the same, since Prototype does not update the feature extractor of the model. FR (Eastwood et al., 2022) and VM are not plotted in (a) because they had huge errors.
Figure 2 plots the reconstruction error $L$ versus $n$ on the three datasets. The error decreased as $n$ increased for all methods, but SSA reduced the error with a smaller $n$ than in those of the baselines, indicating that it could make the target features fit within the source subspace. Especially in the case of Biwi Kinect (c), the baseline methods produced larger errors than Source; i.e., they broke the learned subspace.
In terms of feature alignment, Figure 4 in Section D.2 visualizes and evaluates the feature gap between the domains in the subspace.
Another effect of subspace detection. Subspace detection has another effect on feature alignment. We used the KL divergence between two diagonal Gaussian distributions in Equation 3 to measure the distribution gap between the source and target features, under the assumption that the features follow a Gaussian distribution. Here, while it is not clear that the assumption actually holds especially when features are output through activation functions like ReLU as in ResNet, we argue that the subspace detection of SSA has, in addition to an effective feature alignment, the effect of making such features follow a distribution close to a Gaussian.
@@
Figure 3: Histograms of three randomly selected target feature dimensions. Left: Original features. Right: Projected features.
We visualized the histograms of the target features $\mathbf{z}_{i}^{\text{t}}$ extracted with the source model and ones projected to the source subspace with Equation 6. Figure 3 show the histograms of three randomly selected dimensions of the original and projected features of the three datasets. In the left column of each figure, the histograms of the original features concentrate on zero because of the ReLU activation and do not follow a Gaussian distribution, which makes the KL divergence computation with Equation 3 inaccurate. On the other hand, the histograms of the projected features in the right columns are close to Gaussians. Thus, subspace detection makes it easier to align the features with the Gaussian KL divergence.
The reason why the projected features follow a Gaussian distribution can be interpreted as follows. From Equation 6, the $k$-th element of a projected feature vector $\tilde{z}_{i,k}^{\text{t}}$ is
@@ 5.1 Unsupervised Domain Adaptation
Unsupervised domain adaptation (UDA) has been actively studied as a way to transfer knowledge in the source domain to the target domain (Csurka, 2017). Theoretically, it is known that the upper bound of the error on the target domain includes a distribution gap term between the source and target domains (Ben-David et al., 2010; Ganin et al., 2016; Nguyen et al., 2022). For regression, Cortes & Mohri (2011) theoretically explored regression UDA. RSD (Chen et al., 2021) and DARE-GRAM (Nejjar et al., 2023) take into account that the feature scale matters in regression and explicitly align the feature scale during the feature alignment. However, UDA requires the source and target datasets to be accessed simultaneously during training, which can be restrictive when datasets cannot be accessed due to privacy or security concerns, or storage limitations.
More recently, source-free domain adaptation (SFDA), which does not access the source dataset during adaptation, has been studied. The SFDA setting is similar to TTA in that SFDA adapts models with only unlabeled target data. However, SFDA requires to store the whole target dataset and access the dataset for multiple epochs to train additional models (Li et al., 2020; Xia et al., 2021; Chu et al., 2022; Sanyal et al., 2023; He et al., 2024) or perform clustering (Liang et al., 2020). On the other hand, TTA does not train additional models nor access the target dataset for multiple epochs, which enables instant adaptation with low computational resource and storage.